A BULLET OF MASS 10gm MOVING WITH A VELOCITY OF 400m/s gets embedded in a freely suspended wooden block of mass 900gm . WHAT IS THE VELOCITY ACQUIRED BY THE BLOCK ?


1
Expert's answer
2021-09-19T09:25:31-0400

Write the momentum conservation: initial momentum of the bullet equals the final momentum of the bullet and the block:


"mv=u(M+m)."

According to momentum conservation, the velocity of the block with the bullet embedded is

"u=v\frac{m}{M+m}=4.4\text{ m/s}."